# Noisy Timing Behavior is a Feature of Central Compact Object Pulsars

### v1

@@ Abstract
We present a timing study of the three known central compact object (CCO) pulsars, isolated cooling neutron stars in supernova remnants, using Chandra, XMM-Newton and NICER observations spanning two decades. Relative to canonical young pulsars, CCOs are spinning down at a very slow rate $|\dot{f}|<10^{-15}$ s-2, implying a surface dipole magnetic field strength $B_{s}<10^{11}$ G that is too weak to account for their X-ray emitting hot spots. Two CCO pulsars with sufficiently long monitoring, 1E 1207.4-5209 and PSR J0821-4300, are seen to deviate from steady spin-down; their timing residuals can be modeled by one or more glitches in $f$ and $\dot{f}$, or alternatively by extreme timing noise. For the third CCO pulsar, PSR J1852++0040, the sparse temporal coverage was insufficient to detect such effects. Glitch activity and timing noise in large samples of rotation-powered pulsars correlate best with $\dot{f}$, while the timing irregularities of the first two CCOs are extreme compared to pulsars of the same $\dot{f}$. Nevertheless, timing activity in CCOs may arise from properties that they share with other young but more energetic pulsars: high internal temperature, strong buried magnetic field and superfluid behavior. Alternatively, continuing low-level accretion of supernova debris is not ruled out as a source of timing noise in CCOs.
@@ 1 Introduction
Central compact objects (CCOs) comprise a dozen or so cooling neutron stars (NSs) in supernova remnants (SNRs), defined and distinguished from canonical, young rotation-powered pulsars by their steady surface thermal X-ray emission, lack of surrounding pulsar wind nebula, and non-detection at any other wavelength (Pavlov et al. 2002; see De Luca 2017 for a review). Considering that their faintness and spectral softness are obstacles to widespread discovery in the Galactic plane, CCOs may be as common as more conspicuous NS classes, e.g., magnetars, and may represent a major channel of NS birth.
@@
Pulsations have been detected in only three of the CCOs: 1E 1207.4-5209 in SNR PKS 1209-51/52, PSR J0821-4300 in SNR Puppis A, and PSR J1852++0040 in SNR Kes 79. They have typical periods ($0.424,0.112,0.104$ s) but small spin-down rates $\dot{f}=(-1.2,-6.8,-7.9)\times 10^{-16}$ s-2, implying surface dipole magnetic field strengths of $(9.8,2.9,3.1)\times 10^{10}$ G, respectively, which are exceptionally weak for such young pulsars (Gotthelf et al., 2013; Halpern & Gotthelf, 2010). Furthermore, the surface magnetic field strength of 1E 1207.4-5209 inferred from spin-down, $B_{s}=9.8\times 10^{10}$ G, is close to $B\approx 8\times 10^{10}$ G, the value measured from its unique series of spectroscopic absorption features (Bignami et al., 2003), interpreted as the electron cyclotron fundamental at 0.7 keV and its harmonics. The similar values of field strength obtained by the two independent methods is convincing evidence of the weak surface $B$ field for at least this one CCO.
@@
The remainder of the CCOs, while having similar continuum spectral properties as the CCO pulsars, have eluded searches for pulsations (Alford & Halpern, 2023). It is postulated that they may have even weaker magnetic fields, more uniform surface temperature distributions, or an unfavorable viewing geometry. In fact, it is difficult to create the hot spots needed to produce the X-ray pulsations observed in CCOs assuming only their weak surface dipoles. The only mechanism thought to be capable of creating a nonuniform surface temperature is anisotropic heat conduction in a strong magnetic field. As recounted in Gotthelf & Halpern (2018), models that invoke strong non-dipolar fields in the crust have long been studied. More recently, e.g., Igoshev et al. (2021) calculated the magnetothermal evolution of strong crustal fields generated by a stochastic dynamo, finding a tangled geometry with negligible dipole component. This creates factors of  2 difference in surface temperature on small scales, sufficient to produce the observed pulsations in CCOs.
@@
Postulating the existence of strong non-dipole crustal fields was also the key to an explanation for the conspicuous absence of CCO descendants, the large population of weak $B$-field pulsars that should remain in the same location of the $P-\dot{P}$ diagram after their SNRs have faded. If CCOs are actually born with a canonical NS magnetic dipole field that was promptly buried by fallback of a small amount $(\sim 10^{-4}\,M_{\odot})$ of supernova debris, the buried field will diffuse back to the surface on a timescale of $\sim 10^{5}$ years (Muslimov & Page, 1995; Bernal et al., 2010; Ho, 2011; Viganò & Pons, 2012). A CCO will move vertically up in the $P-\dot{P}$ diagram as its dipole field grows back (Bogdanov et al., 2014; Ho, 2015; Luo et al., 2015), merging with the population of ordinary radio pulsars.
@@
The discovery of a glitch and/or strong timing noise from 1E 1207.4-5209 (Gotthelf & Halpern, 2018, 2020) was of timely relevance to these theories of CCOs. Glitches are thought to result from either “starquakes”, stress relief of the NS crust, or from unpinning of superfluid vortices in the inner part of the NS crust (e.g., Link et al., 1998). A review of glitch observations and theory is given by Antonopoulou et al. (2022), while a comprehensive compilation of glitch statistics can be found in Espinoza et al. (2011) and Fuentes et al. (2017). Timing noise has been attributed to variability in the interaction of the crustal superfluid with the Coulomb lattice of the solid crust (Jones, 1990), turbulence of the superfluid (Melatos & Link, 2014), or fluctuations in the structure of the magnetosphere, e.g., state switching (Lyne et al., 2010). Analysis methods for timing noise are reviewed in Parthasarathy et al. (2019) and Namkham et al. (2019).
@@
The timing irregularities in 1E 1207.4-5209 were extreme relative to the general pulsar population, in which glitch activity is correlated primarily with frequency derivative $\dot{f}$. Naively, glitches are unexpected in CCOs because of their small $\dot{f}$. However, the idea of stronger internal $B$ fields, needed to explain CCO surface hot spots and population statistics, may have an additional application in producing the timing irregularities of CCOs. Ho (2015) proposed that glitch activity could be triggered by the motion of strong magnetic fields through the NS crust, interacting with the neutron superfluid there, and that glitches would be a way of identifying CCO descendants.
@@
To add complexity, it has not been ruled out that an entirely different process, low-level accretion in the propeller regime, could cause the observed timing noise, specifically the observed changes in $\dot{f}$ (Halpern et al., 2007; Gotthelf & Halpern, 2018). This, because the required accretion rate of $\leq 10^{-11}$ g s-1 makes a negligible contribution to the observed luminosity, while the mass supply could be a residual fallback disk of $\sim 10^{-6}\,M_{\odot}$, much less mass than may have fallen back on the NS initially. The spin parameters of CCOs fall in regime where dipole braking and accretion torque noise could be comparable.
@@
The subject of this paper is the long-term spin-down evolution of all three CCO pulsars spanning the two decades since their discovery. In Section 2 we present new Chandra, XMM-Newton, and NICER timing observations which we use to update timing solutions, explore alternative timing models, and identify timing irregularites in PSR J0821-4300 for the first time. Section 3 discusses the implications of the candidate glitches and timing noise, and compares the CCO results with those of the general pulsar population. Section 4 summarizes our conclusions.
2 Timing Analysis
@@
We obtained new timing observations for the three CCO pulsars as part of our continued monitoring programs using the Chandra and XMM-Newton observatories. We were able to achieve better temporal coverage by requesting observations from the two missions, allowing for their unique scheduling constraints. With the onset of the NICER mission, we continued our coverage of 1E 1207.4-5209 using NICER exclusively. The previously published data sets also included here are fully described in our earlier work (1E 1207.4-5209 — Gotthelf & Halpern, 2007; Halpern & Gotthelf, 2011; Gotthelf et al., 2013; Halpern & Gotthelf, 2015; Gotthelf & Halpern, 2018, 2020), (PSR J0821-4300 — Gotthelf & Halpern, 2009; Gotthelf et al., 2013), and (PSR J1852++0040 — Gotthelf et al., 2005; Halpern et al., 2007; Halpern & Gotthelf, 2010).
@@
All data sets were reprocessed and reduced using the latest software for each mission, following the methods described in our earlier papers. Photon arrival times were converted to barycentric dynamical time (TDB) using the DE405 solar system ephemeris and the Chandra coordinates given in the tables below for each object. For PSR J0821-4300, we account for its high proper motion using the ephemeris of Mayer et al. (2020). Photons were extracted using a circular aperture optimized for each target and instrument, based on the local SNR contamination in the aperture.
@@
For each observation, we generated a phase zero time-of-arrival (ToA) measurement to use in determining the long-term spin evolution of the NS, as follows. Starting from the filtered event files, we measured the pulse period at each epoch using a periodogram search for the maximum power around the expected frequency, as informed by our earlier work. For 1E 1207.4-5209 and PSR J0821-4300, their pulse profiles are well-characterized by a sine function, and we computed phase zero from the Fourier coefficients of the unbinned photon arrival times. We define phase zero as the peak of the sine function (see Gotthelf & Halpern, 2020, for details). For PSR J1852++0040, which has a much broader pulse profile, a high-significance pulse profile template was generated by adding all the folded profiles, aligned by cross-correlating with an iterated template. We define phase zero for this CCO as the middle of the valley in the pulse profile.
@@
Utilizing the TEMPO software (Hobbs et al., 2006), we initiated a phase-connected timing solution to a subset of ToAs for each pulsar, using a model for the rotation phase of the pulsar that includes one or two of its frequency derivatives:
@@
Generally, we obtained optimally spaced sets of observations that allow us to start and to maintain a phase-connected timing solution that accounts for every turn of the pulsar, with increased accuracy, by bootstrapping to longer intervals. This is usually possible unless the interval to the next ToA is too large to predict its cycle count by less that $\sim 10\%$ in phase. As described below, due to data gaps and timing irregularities, it was necessary at some point to restart a new solution for each pulsar to continue to follow and/or interpret its spin-down behavior. However, if the timing residuals begin to deviate significantly and systematically from the trial solution, it signifies that the ephemeris no longer predicts the pulsar’s spin-down, and more complex models need to be considered.
2.1 1E 1207.4-5209

### v2

@@ Abstract
We present a timing study of the three known central compact object (CCO) pulsars, isolated cooling neutron stars in supernova remnants (SNRs), using observations from Chandra, XMM-Newton and NICER spanning two decades. Relative to canonical young pulsars, CCOs are spinning down at a very slow rate $|\dot{f}|<10^{-15}$ s-2, implying a surface dipole magnetic field strength $B_{s}<10^{11}$ G that is too weak to account for their X-ray-emitting hot spots. Two CCO pulsars with sufficiently long monitoring, 1E 1207.4-5209 and PSR J0821-4300, are seen to deviate from steady spin-down; their timing residuals can be modeled by one or more glitches in $f$ and $\dot{f}$, or alternatively by extreme timing noise. For the third CCO pulsar, PSR J1852++0040, the sparse temporal coverage was insufficient to detect such effects. Glitch activity and timing noise in large samples of rotation-powered pulsars correlate best with $\dot{f}$, while the timing irregularities of the first two CCOs are extreme compared to pulsars of the same $\dot{f}$. Nevertheless, timing activity in CCOs may arise from properties that they share with other young but more energetic pulsars: high internal temperature, strong buried magnetic field, and superfluid behavior. Alternatively, continuing low-level accretion of supernova debris is not ruled out as a source of timing noise in CCOs.
@@ I Introduction
Central compact objects (CCOs), comprising a dozen or so cooling neutron stars (NSs) in supernova remnants (SNRs), are defined and distinguished from canonical, young rotation-powered pulsars by their steady surface thermal X-ray emission, lack of surrounding pulsar wind nebula, and nondetection at any other wavelength [^1] Recently, the first detection of weak radio pulsations from a CCO, 1E 1207.4-5209, was reported (Zhang et al., 2025). (Pavlov et al. 2002; see De Luca 2017 for a review). Considering that their faintness and spectral softness are obstacles to widespread discovery in the Galactic plane, CCOs may be as common as more conspicuous NS classes, e.g., magnetars, and may represent a major channel of NS birth.
@@
X-ray pulsations, evidently due to nonuniform surface temperature, have been detected from only three of the CCOs: 1E 1207.4-5209 in SNR PKS 1209-51/52, PSR J0821-4300 in SNR Puppis A and PSR J1852++0040 in SNR Kes 79. They have typical spin periods $P=(0.424,0.112,0.104$) s but small spin-down rates $\dot{f}=(-1.2,-6.8,-7.9)\times 10^{-16}$ s-2, implying surface dipole magnetic field strengths of $(9.8,2.9,3.1)\times 10^{10}$ G, respectively, which are exceptionally weak for such young pulsars (Gotthelf et al., 2013; Halpern & Gotthelf, 2010).
The ages of the pulsars are the same as those of their SNRs, estimated as $t=(14,10,5)\times 10^{3}$ yr for PKS 1209-51/52 (Eppens et al., 2024), Puppis A (Aruga et al., 2022) and Kes 79 (Mayer & Becker, 2021), respectively. In contrast, when a pulsar “characteristic age” is quoted, defined as $\tau_{c}\equiv P/2\dot{P}$, the value is in the range $(2-3)\times 10^{8}$ yr for CCOs. Because the relation between $\tau_{c}$ and true age $t$ is $t=\tau_{c}[1-(P_{0}/P)^{2}]$ for dipole braking with constant magnetic field, this spectacular difference simply indicates that the current spin period $P$ of the pulsar is likely identical to its otherwise unknown period at birth $P_{0}$ because its $B$ field is so weak.
Crucially, the surface magnetic field strength of 1E 1207.4-5209 inferred from spin-down, $B_{s}=9.8\times 10^{10}$ G, is close to $B\approx 8\times 10^{10}$ G, the value measured from its unique series of spectroscopic absorption features (Bignami et al., 2003) interpreted as the electron cyclotron fundamental at 0.7 keV and its harmonics. That similar values of field strength are obtained by the two independent methods is persuasive evidence of the weak surface $B$ field for at least this CCO.
The remainder of the CCOs, having similar continuum spectral properties, have eluded searches for X-ray pulsations (Alford & Halpern, 2023). It is postulated that they may have even weaker magnetic fields, more uniform surface temperature distributions, or unfavorable geometry of hot spots. Alford et al. (2022) derived that the hot spots on the surface of PSR J0821-4300 are probably close to its spin axis, and this could be the case for CCOs in general (Alford & Halpern, 2023), which would limit the possible amplitude and detectability of pulsation. A similar conclusion about the magnetic axis of 1E 1207.4-5209, that it is nearly aligned with the spin, was tentatively drawn from the polarization swing and width of its radio pulse (Zhang et al., 2025).
@@
It is even difficult to create the hot spots needed to produce X-ray pulsations in CCOs assuming only their weak surface dipoles. The only mechanism thought to be capable of creating a nonuniform surface temperature is anisotropic heat conduction in a strong magnetic field. As recounted in Gotthelf & Halpern (2018), models that include strong nondipolar components, mainly toroidal ones in the crust, have long been studied. More recently, Igoshev et al. (2021) calculated the magnetothermal evolution of strong crustal fields generated by a stochastic dynamo, finding a tangled geometry with negligible dipole component. This creates factors of $\sim 2$ difference in surface temperature on small scales, sufficient to produce the observed pulsations in CCOs.
@@
Postulating a temporary suppression of the surface dipole field was the key to an explanation for the conspicuous absence of CCO descendants, the large population of weak $B$-field pulsars that should remain in the same location of the $P-\dot{P}$ diagram after their SNRs have faded. If CCOs are actually born with a canonical NS magnetic dipole field that was promptly buried by fallback of a small amount $(\sim 10^{-4}\,M_{\odot})$ of supernova debris, the buried field will diffuse back to the surface on a timescale of $\sim 10^{5}$ yr (Muslimov & Page, 1995; Bernal et al., 2010; Ho, 2011; Viganò & Pons, 2012). A CCO will move vertically up in the $P-\dot{P}$ diagram as its dipole field grows back (Bogdanov et al., 2014; Ho, 2015; Luo et al., 2015), merging with the population of ordinary radio pulsars.
@@
The discovery of a glitch and/or strong timing noise from 1E 1207.4-5209 (Gotthelf & Halpern, 2018, 2020) was of timely relevance to these theories of CCOs. Glitches are thought to result from “starquakes”, stress relief of the NS crust, or from unpinning of superfluid vortices in the inner part of the NS crust (e.g., Link et al., 1998). A review of glitch observations and theory is given by Antonopoulou et al. (2022), while comprehensive compilations of glitch statistics can be found in Espinoza et al. (2011), Fuentes et al. (2017) and Basu et al. (2022). Timing noise has been attributed to variability in the interaction of the crustal superfluid with the Coulomb lattice of the solid crust (Jones, 1990), turbulence of the superfluid (Melatos & Link, 2014), or fluctuations in the structure of the magnetosphere, e.g., state switching (Lyne et al., 2010). Analysis methods for timing noise are reviewed in Parthasarathy et al. (2019) and Namkham et al. (2019).
@@
The timing irregularities in 1E 1207.4-5209 were extreme relative to the general pulsar population, in which glitch activity is mainly correlated with the frequency derivative $\dot{f}$. Glitches are therefore unexpected in CCOs due to their small $\dot{f}$. However, the idea of stronger internal $B$ fields, needed to explain the CCOs’ surface hot spots and population statistics, may find additional application in producing the timing irregularities of CCOs. Ho (2015) proposed that glitches could be triggered by the motion of strong magnetic fields through the NS crust, interacting with the neutron superfluid there, and that glitches would be a way of identifying CCO descendants.
@@
To add complexity, it has not been ruled out that an entirely different process, low-level accretion in the propeller regime, could cause the observed timing noise, specifically the observed changes in $\dot{f}$ (Halpern et al., 2007; Gotthelf & Halpern, 2018). This is because the required accretion rate of $\leq 10^{11}$ g s-1 makes a negligible contribution to the observed luminosity, while the mass supply could be a residual fallback disk of $\sim 10^{-6}\,M_{\odot}$, much less mass than may have fallen back on the NS initially. The spin parameters of CCOs fall into a regime where dipole braking and accretion torque noise could be comparable.
@@
The subject of this paper is the long-term spin-down evolution of all three CCO pulsars spanning the two decades since their discovery. In Section II we present new observations from Chandra, XMM-Newton, and NICER that we use to update timing solutions, explore alternative timing models, and identify timing irregularities in PSR J0821-4300 for the first time. Section III discusses the implications of the candidate glitches and timing noise and compares the CCO results with those of the general pulsar population. Section IV summarizes our conclusions.
II Timing Analysis
@@
We obtained new timing observations for the three CCO pulsars as part of our continued monitoring programs using the Chandra and XMM-Newton observatories. We were able to achieve better temporal coverage by requesting observations from the two missions, allowing for their unique scheduling constraints. With the onset of the NICER mission, we continued our coverage of 1E 1207.4-5209 using NICER exclusively. The previously published datasets also included here are fully described in our previous work (1E 1207.4-5209 — Gotthelf & Halpern, 2007; Halpern & Gotthelf, 2011; Gotthelf et al., 2013; Halpern & Gotthelf, 2015; Gotthelf & Halpern, 2018, 2020), (PSR J0821-4300 — Gotthelf & Halpern, 2009; Gotthelf et al., 2013), and (PSR J1852++0040 — Gotthelf et al., 2005; Halpern et al., 2007; Halpern & Gotthelf, 2010). All of the observations used here are listed in the Appendix.
@@
All datasets were reprocessed and reduced using the latest software for each mission, following the methods described in our previous papers. Photon arrival times were converted to barycentric dynamical time (TDB) using the DE405 solar system ephemeris and the Chandra coordinates given in the tables below for each object. For PSR J0821-4300, we account for its high proper motion using the ephemeris of Mayer et al. (2020). Photons were extracted using a circular aperture optimized for each target and instrument, based on the local SNR contamination in the aperture.
@@
For each observation, we generated a phase zero time-of-arrival (ToA) measurement to use in determining the long-term spin evolution of the NS, as follows. Starting from the filtered event files, we measured the pulse period at each epoch using a periodogram search for the maximum power around the expected frequency, as informed by our earlier work. For 1E 1207.4-5209 and PSR J0821-4300, their pulse profiles are well characterized by a sinusoidal function, and we computed phase zero from the Fourier coefficients of the unbinned photon arrival times. For consistency with Gotthelf & Halpern (2020), we define phase zero as the minimum of the pulse for 1E 1207.4-5209. For PSR J0821-4300 we use the maximum of the pulse as phase zero. For PSR J1852++0040, which has a much broader pulse profile, a high-significance pulse profile template was generated by adding all the folded profiles, aligned by cross-correlating with an iterated template. We define phase zero for this CCO as the middle of the valley in the pulse profile.
@@
Using the TEMPO software (Nice et al., 2015), we initiated a phase-connected timing solution to a subset of ToAs for each pulsar, using a model for the rotation phase of the pulsar that includes one or two of its frequency derivatives:
@@
Generally, we obtained optimally spaced sets of observations that allow us to start and maintain a phase-connected timing solution that accounts for every turn of the pulsar, with increased accuracy, by bootstrapping to longer intervals. This is usually possible unless the interval to the next ToA is too large to predict its cycle count by less than $\sim 10\%$ in phase. As described below, due to data gaps and timing irregularities, it was necessary at some point to restart a new solution for each pulsar to continue to follow and/or interpret its spin-down behavior. However, if the timing residuals begin to deviate significantly and systematically from the trial solution, it signifies that the ephemeris no longer predicts the pulsar’s spin-down, and more complex models need to be considered.
II.1 1E 1207.4-5209 in PKS 1209-51/52